OpenStudy (anonymous):

does anyone know how to solve this ? given an equation : y(t) = 5x^2 (t) - 4 dy/dt a. proof that the system is linear or nonlinear b. proof that the system is time invariant or time variant

OpenStudy (anonymous):

y=5x^2-4dy/dt 4dy/dt=5x^2-y dy/dt=5/4x^2-1/4y dy=5/4x^2dt-1/4ydt integrate the equation y=int(5/4x^2dt)-1/4int(ydt) its nonlinear

OpenStudy (anonymous):

hey its complex we are provided with y(t) in right hand side of teh equation not axactly y.. @sara17

OpenStudy (anonymous):

it will be solved if we apply laplace and inverse laplace tranform on the above equation..